OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8

OpenStudy (anonymous):

1st eq:3x + y = 6 2(3x+y)=2(6) 6x+2=12 2nd eq:6x + 2y = 8 as 1st eq and 2nd eq has a different solution so it has no answers